Using the transformation T=z^2/2, reduce the equation f(x,y,z,p,q)= x(y^2+z^2×q^2)- zy^2×p=0 to a form f(P,x)= g(Q,y) where p= ∂T/∂x, Q= ∂T/∂y
1
Expert's answer
2020-11-09T20:06:03-0500

Given equation is,

"f(x,y,z,p,q) = x(y^2+z^2\u00d7q^2)- zy^2p= xy^2+xz^2q^2- zy^2p"


Since, Transformation is, "T = \frac{z^2}{2}"


Now,

"z = \sqrt{2T}"

"p =\frac{\partial z}{\partial x} = \sqrt{2}\frac{1}{2} T^{-1/2}\frac{\partial T}{\partial x} =\frac{1}{\sqrt{2T}} \frac{\partial T}{\partial x} =\frac{1}{\sqrt{2T}} P"


"q=\frac{\partial z}{\partial y} = \sqrt{2}\frac{1}{2} T^{-1/2}\frac{\partial T}{\partial y} =\frac{1}{\sqrt{2T}} \frac{\partial T}{\partial y} =\frac{1}{\sqrt{2T}} Q"


Putting value of z, p and q in the given equation,

"xy^2+x(2T)(\frac{1}{\sqrt{2T}} Q)^2 -\sqrt{2T} y^2 (\frac{1}{\sqrt{2T}} P)=0"


"xy^2+x(2T)(\frac{1}{{2T}} Q^2) - y^2 ( P)=0"

"xy^2+xQ^2 - y^2 P=0"

"y^2(x-P) = -xQ^2"


"\frac{y^2}{Q^2} = \frac{x}{(P-x)}"

It can written as,

"g(Q,y) = f(P,x)"

which is the required result.